Daily Updates - April 27, 2005
Opportunity Status at end of sol 438-442
Opportunity keeps driving southward and studying new locations despite a disabled right-front steering motor. Opportunity has driven about 110 meters (361 feet) without use of that motor. The miniature thermal emission spectrometer began returning good data again. That instrument was in use when the rover stopped operating for a software reset on sol 440. The rover continues making scientific observations while engineers diagnose the cause of the reset.
Sol 438 (April 17, 2005):
Opportunity made remote-sensing observations.
Opportunity performed a 13-filter panoramic camera observation to study soil in a trench that was scooped by a wheel when the rover turned to a good communications orientation after its sol 437 drive. Opportunity followed the camera observations with an 80-meter (262-foot) drive south.
The team's plan was for Opportunity to make remote-sensing observations and then drive farther south. Panoramic camera imaging and some miniature thermal emission spectrometer observations were successfully completed. A miniature thermal emission spectrometer observation was underway when a software reset occurred at approximately 12:45 Mars local solar time.
The team prepared a recovery plan responding to the software reset the sol before. The plan included transmission of data acquired prior to and during the sol 440 event. Some of this data was returned during a downlink through the Odyssey orbiter on sol 441. Additional data were requested for transmission on sol 442 in hopes of pinpointing the cause of the software reset. Opportunity is otherwise healthy.
The team told Opportunity to perform remote science and study the surface at its present location while the engineering evaluation continued.
As of sol 442 (April 22, 2005) Opportunity's odometry total is 5,306 meters (3.30 miles).
Opportunity Daily Update Archive